JOINT DIFFERENTIAL GEOMETRY/PDE & INVERSE PROBLEMS & RAINWATER SEMINAR

Tuesday, November 23, 2004 Padelford C-401 2:30 PM (Note: special day, time and room!)

Boundary rigidity and volume minimality for almost flat metrics

Dmitri Burago (Pennsylvania State University)

A compact Riemannian manifold with boundary is said to be boundary rigid if its metric is uniquely determined (up to an isometry) by the distances between the boundary points; and is said to be a minimal filling if it has the least volume among all compact Riemannian manifolds with the same boundary and the same or greater boundary distances. I will discuss the following result of a recent joint work with S. Ivanov: Euclidean regions with Riemannian metrics sufficiently close to a Euclidean one are minimal fillings and boundary rigid.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.